OpenStudy (anonymous):

Show that A = [(3, 2, 4), (2, 0, 2), (4, 2, 3)] a 3x3 matrix, is diagnolizable even though one eigenvector has algebraic multiplicity 2. Do this by brute force computation. Why would you expect this to be true even without calculation?

OpenStudy (ash2326):

@lgbasallote @UnkleRhaukus @Ishaan94 Please help here

OpenStudy (anonymous):

I am sorry I don't know what eigenvector is

OpenStudy (unklerhaukus):

\[\textbf A=\left(\begin{array} {ccc} 3&2&4\\2&0&2\\4&2&3 \end{array}\right)\left(\begin{array}{ccc}x_1\\x_2\\x_3\end{array}\right) \]

OpenStudy (unklerhaukus):

i guess we should find the eigen vectors,

OpenStudy (unklerhaukus):

\[\textbf A\textbf x=\lambda \textbf x\]\[(\textbf A-\textbf I\lambda)\textbf x=0\]\[\left|\textbf A-\textbf I\lambda\right|=0\]\[\left|\begin{array} {ccc} 3-\lambda&2&4\\2&0-\lambda&2\\4&2&3-\lambda \end{array}\right|=0\]

OpenStudy (unklerhaukus):

\[(3-\lambda)\left[(-\lambda)\times(3-\lambda)-2\times2\right]-2[2\times(3-\lambda)-2\times4]+4[2\times2-(-\lambda\times4)]=0\] \[(3-\lambda)\left[-3\lambda+\lambda^2-4\right]-2\left[6-2\lambda-8\right]+4\left[4+4\lambda\right]=0\]

OpenStudy (unklerhaukus):

\[-9\lambda+3\lambda^2-12-3\lambda^2-\lambda^3+4\lambda-12+4\lambda+16+16+16\lambda=0\] \[-\lambda-\lambda^3+16\lambda+8=0\]

OpenStudy (unklerhaukus):

which dosent factor nicely, so i must have made a mistake somewhere ..

OpenStudy (zarkon):

your \(-3\lambda^2\) should be \(+3\lambda^2\)

OpenStudy (zarkon):

giving you \(6\lambda^2\)

OpenStudy (zarkon):

\[-9\lambda+3\lambda^2-12+3\lambda^2+\lambda^3+4\lambda-12+4\lambda+16+16+16\lambda=0\]